OpenStudy (anonymous):

what is the sum of these mixed numbers ? 3 1/3 + 2 1/6

OpenStudy (anonymous):

5 3/6

OpenStudy (anonymous):

3+(1/3) becomes 10/3 2+(1/6) becomes 13/6 10/3 then becomes 20/6 20/6+13/6=33/6 you can then divide by 3 to get 11/2

OpenStudy (anonymous):

5 3/6

OpenStudy (anonymous):

not 5 5/6 ?

OpenStudy (anonymous):

no 3 1/3 = 3 2/6 2 1/6 +3 2/6 = 5 3/6

OpenStudy (anonymous):

my options are 5 5/6 ,5 1/2 , 5 2/3

OpenStudy (anonymous):

5 1/2

OpenStudy (anonymous):

how did you get 5+5/6? 11/2 = 5+(3/6) or 5+(1/2) 5=10/2 (3/6)=(1/2) therefore, your correct answer is 5+(1/2)

OpenStudy (anonymous):

Then 5 3/6 would become 5 1/2, sorry I forgot to simplify

OpenStudy (anonymous):

ok i was thinking i needed to simplify